# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ Abstract
Special approximation technique for analysis of different characteristics of states of multipartite infinite-dimensional quantum systems is proposed and applied to the study of the relative entropy of $\pi$-entanglement and its regularisation.
In particular, by using this technique we obtain simple sufficient conditions for local continuity (convergence) of the regularized relative entropy of $\pi$-entanglement.
We establish a finite-dimensional approximation property for the relative entropy of entanglement and its regularization that allows us to generalize to the infinite-dimensional case the results proved in the finite-dimensional settings.
We also show that for any multipartite state with finite energy the infimum in the definition of the relative entropy of $\pi$-entanglement can be taken over the set of finitely-decomposable $\pi$-separable states with finite energy.
@@ 1 Introduction
The FA-property of a state $\rho$ implies the finiteness of its entropy [32], but the converse implication is not true (this is shown in [2] by constructing a subtle example of a state with finite entropy not possessing the FA-property). In Section 2.2 the following simple sufficient condition for the FA-property is proved: the FA-property holds for a state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}$ provided that [^2] This condition strengthens the condition obtained in [32] which states that the FA-property holds for a state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}$ provided that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{q}i<+\infty$ for some $q>2$.
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}i<+\infty.$
This condition shows that the FA-property holds for all states whose eigenvalues tend to zero faster than $[i\ln^{q}i]^{-1}$ as $i\to+\infty$ for some $q>3$, in particular, it holds for all Gaussian states playing essential role in quantum information theory.
Many characteristics of a state of a $n$-partite quantum system $A_{1}...A_{n}$ have the form of a function
$f(\rho\,|\,p_{1},...,p_{l})$
@@
on the set $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ of states of this system depending on some parameters $p_{1},...,p_{l}$ (other states, quantum channels, quantum measurements, etc.). If $\rho$ is a state of infinite-dimensional $n$-partite quantum system $A_{1}...A_{n}$ then we may approximate it by the sequence of states
$\rho_{r}=Q_{r}\rho\hskip 1.0ptQ_{r}[\mathrm{Tr}Q_{r}\rho\hskip 1.0pt]^{-1},\quad Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n},$ (1)
@@
where $P_{r}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account). Naturally, the question arises under what conditions
$f(\rho_{r}\,|\,p_{1},...,p_{l})\quad\textrm{tends to}\quad f(\rho\,|\,p_{1},...,p_{l})\quad\textrm{as}\;\;r\to\infty$
@@
uniformly on $p_{1},...,p_{l}$. Our main technical result asserts that for given $m\leq n$ the FA-property of the marginal states $\rho_{A_{1}}$,…, $\rho_{A_{m}}$ of a state $\rho$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ guarantees this uniform convergence for a wide class of functions $f$ (depending on $m$). In fact, the more general assertion is valid in which the states $\rho_{r}$ are defined by formula (1) via the operators $Q_{r}=P_{r}^{s_{1}}\otimes...\otimes P_{r}^{s_{l}}\otimes I_{R}$, where $\{s_{1},...,s_{l}\}$ is any subset of $\{1,...,n\}$ and $R=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{l}}$. For a smaller class of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ the above result is valid with the FA-property of the states $\rho_{A_{1}}$,…, $\rho_{A_{m}}$ replaced by the weaker condition of finiteness of their entropy (Theorem 3 in Section 3).
@@
The proposed approximation technique allows us to establish the lower semicontinuity of the regularized relative entropy of $\pi$-entanglement $E^{\infty}_{R,\pi}$ on the subset $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of states $\rho$ that have at least $n-1$ marginal states with the FA-property. Moreover, it is proved that
$\liminf_{k\to+\infty}E^{\infty}_{R,\pi}(\rho_{k})\geq E^{\infty}_{R,\pi}(\rho_{0}),$
for arbitrary sequence $\{\rho_{k}\}$ converging to a state $\rho_{0}$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
By combining this observation with the results from [15] we obtain simple sufficient conditions for local continuity of the regularized relative entropy of $\pi$-entanglement $E^{\infty}_{R,\pi}$, i.e. conditions under which the limit relation
$\lim_{k\to+\infty}E^{\infty}_{R,\pi}(\rho_{k})=E^{\infty}_{R,\pi}(\rho_{0})<+\infty$
holds for a given sequence $\{\rho_{k}\}$ converging to a state $\rho_{0}$.
The approximation technique is used also to obtain several observations concerning definition of the relative entropy of $\pi$-entanglement. In particular, we prove that for all states in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ this quantity can be defined as the relative entropy distance to the set of all finitely decomposable separable states (despite the existence of finitely-non-decomposable and countably-non-decomposable separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$).
We establish a finite-dimensional approximation property for the relative entropy of entanglement and its regularization in $n$-partite quantum systems that allows us to generalize to the infinite-dimensional case the results proved in the finite-dimensional settings. Examples of using this property are presented.
Finally, we consider energy-constrained versions of the $n$-partite relative entropy of $\pi$-entanglement. It is proved, in particular, that for any state in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ with finite energy the infimum in the definition of the relative entropy of $\pi$-entanglement can be taken over all finitely-decomposable separable states with finite energy (provided that the Hamiltonians of individual subsystems satisfy a certain condition).
@@ 2.1 Basic notation
The quantum relative entropy for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H})$ can be defined as
$D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)=\sum_{i}\langle\varphi_{i}|\,\rho\ln\rho-\rho\ln\sigma\,|\varphi_{i}\rangle,$
where $\{\varphi_{i}\}$ is the orthonormal basis of eigenvectors of the state $\rho$ and it is assumed that $D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)=+\infty$ if $\,\mathrm{supp}\rho\hskip 1.0pt$ is not contained in $\hskip 1.0pt\mathrm{supp}\hskip 1.0pt\sigma$ [16, 9, 39]. [^3] The support $\mathrm{supp}\rho$ of a positive trace class operator $\rho$ is the closed subspace spanned by the eigenvectors of $\rho$ corresponding to its positive eigenvalues.
@@
The quantum mutual information (QMI) of a state $\,\rho\,$ of a multipartite quantum system $A_{1}\ldots A_{n}$ is defined as (cf.[17, 8, 40])
$I(A_{1}\!:\ldots:\!A_{n})_{\rho}\doteq D(\rho\,\|\,\rho_{A_{1}}\otimes\cdots\otimes\rho_{A_{n}})=\sum_{s=1}^{n}S(\rho_{A_{s}})-S(\rho),$
@@
For any positive (semi-definite) densely defined operator [^4] We assume that a positive operator is a self-adjoint operator [24]. $H$ on a Hilbert space $\mathcal{H}$ and any positive operator $\rho$ in $\mathfrak{T}(\mathcal{H})$ we assume that
$\mathrm{Tr}H\rho=\sup_{n}\mathrm{Tr}P_{n}H\rho\leq+\infty,$ (5)
where $P_{n}$ is the spectral projector of $H$ corresponding to the interval $[0,n]$. Then
@@
is a closed convex nonempty subset of $\mathfrak{S}(\mathcal{H})$ for any $E$ exceeding the infimum of the spectrum of $H$. If $H$ is treated as Hamiltonian of a quantum system associated with the space $\mathcal{H}$ then $\mathfrak{C}_{H,E}$ is the set of states with the mean energy not exceeding $E$.
We will pay a special attention to the class of unbounded densely defined positive operators on $\mathcal{H}$ having discrete spectrum. In Dirac’s notations any such operator $H$ can be represented as follows
@@
on the domain $\mathcal{D}(H)=\{\varphi\in\mathcal{H}\,|\,\sum_{i=1}^{+\infty}h^{2}_{i}|\langle\tau_{i}|\varphi\rangle|^{2}<+\infty\}$, where $\left\{\tau_{i}\right\}_{i=1}^{+\infty}$ is the orthonormal basis of eigenvectors of $H$ corresponding to the nondecreasing sequence $\left\{\smash{h_{i}}\right\}_{i=1}^{+\infty}$ of eigenvalues tending to +$\infty$+ ∞.
@@
and that the maximal value of the entropy on this set is achieved at the Gibbs state $\gamma_{H}(E)\doteq e^{-\beta(E)H}/\mathrm{Tr}e^{-\beta(E)H}$, where the parameter $\beta(E)$ is determined by the equality $\mathrm{Tr}He^{-\beta(E)H}=E\mathrm{Tr}e^{-\beta(E)H}$ [38]. Condition (8) implies that $H$ is an unbounded operator having discrete spectrum, i.e. it has the form (7). So, by the Lemma in [10] the set $\mathfrak{C}_{H,E}$ defined in (6) is compact for any $E$.
@@
If equivalent properties $(i)\textrm{-}(iii)$ hold then the positive operator $H$ in $(ii)$ can be taken diagonizable in any basis of eigenvectors of the state $\rho$.
Proof. The implication $\rm(ii)\Rightarrow(i)$ follows from the results mentioned before. To prove $\rm(i)\Rightarrow(iii)$ it suffices to note that $\lambda^{\rho}_{i}\leq 1/i$ for all $i$.
Assume that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln i<+\infty$. Then it is easy to show the existence of an increasing sequence $\{c_{i}\}_{i=1}^{+\infty}$ such that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}c_{i}\ln i<+\infty$. Then $\rm(ii)$ holds with the operator
@@
where $\{\varphi_{i}\}_{i=1}^{+\infty}$ is any basis of eigenvectors of $\rho$ corresponding to the sequence $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$. $\Box$
@@
which is stronger than condition (8). The condition (10) is valid if $\;\liminf\limits_{i\rightarrow+\infty}h_{i}/\ln^{p}i>0\,$ for some $\,p>2$, where $\{h_{i}\}$ is the sequence of eigenvalues of $H$ [28, Proposition 1]. By Lemma 1 in [28] condition (10) holds if and only if
@@
$F_{H^{2}}(E)\leq F_{H}(\sqrt{E})\quad\forall E\geq E^{2}_{0},$
where $E_{0}$ is the minimal eigenvalue of $H$. This inequality follows from the definitions of $F_{H^{2}}(E)$ and $F_{H}(\sqrt{E})$ due to the inequality
@@
valid for any state $\rho$ in $\mathfrak{S}(\mathcal{H})$ (with possible value +$\infty$+ ∞ in one or both sides). The last inequality is a direct corollary of the concavity of the square root, since

### v2

@@ Abstract
An universal approximation technique for analysis of different characteristics of states of composite infinite-dimensional quantum systems is proposed and used to prove general results concerning the properties of correlation and entanglement measures in such systems. Then these results are applied to the study of three important characteristics: the relative entropy of $\pi$-entanglement, the Rains bound (the unregularized and regularized versions of both characteristics are considered) and the conditional entanglement of mutual information.
In particular, we analyse continuity and convexity properties of the above entanglement measures, prove several results simplifying their definitions and establish a finite-dimensional approximation property for these characteristics that allows us to generalize to the infinite-dimensional case the results proved in the finite-dimensional settings.
@@ 1 Introduction
The FA-property of a state $\rho$ implies the finiteness of its entropy [4], but the converse implication is not true (this is shown in [5] by constructing a subtle example of a state with finite entropy not possessing the FA-property). In Section 2.2 the following simple sufficient condition for the FA-property is proved: the FA-property holds for a state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}$ provided that [^2] This condition strengthens the condition obtained in [4] which states that the FA-property holds for a state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}$ provided that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{q}i<+\infty$ for some $q>2$. $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}i<+\infty$. This condition shows that the FA-property holds for all states whose eigenvalues tend to zero faster than $[i\ln^{q}i]^{-1}$ as $i\to+\infty$ for some $q>3$, in particular, it holds for all Gaussian states playing essential role in quantum information theory.
@@
Many characteristics of a state of a $n$-partite quantum system $A_{1}...A_{n}$ have the form of a function $f(\rho\,|\,p_{1},...,p_{l})$ on the set $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ of states of this system depending on some parameters $p_{1},...,p_{l}$ (other states, quantum channels, quantum measurements, etc.). If $\rho$ is a state of an infinite-dimensional $n$-partite quantum system $A_{1}...A_{n}$ then we may approximate it by the sequence of states
$\rho_{r}=[\mathrm{Tr}Q_{r}\rho\hskip 1.0pt]^{-1}Q_{r}\rho\hskip 1.0ptQ_{r},\quad Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n},$ (1)
@@
where $P_{r}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account). Naturally, the questions arise under what conditions
$f(\rho_{r}\,|\,p_{1},...,p_{l})\quad\textrm{tends to}\quad f(\rho\,|\,p_{1},...,p_{l})\quad\textrm{as}\;\;r\to\infty$ (2)
@@
for given $p_{1},...,p_{l}$ and under what conditions this convergence is uniform on $p_{1},...,p_{l}$. Our main technical result asserts that for given $m\leq n$ the FA-property of the marginal states $\rho_{A_{1}}$,…, $\rho_{A_{m}}$ of a state $\rho$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ implies (2) for a wide class of functions $f$ (depending on $m$). It also gives a sufficient condition under which the convergence in (2) is uniform on $p_{1},...,p_{l}$. In fact, a more general assertion is true in which the states $\rho_{r}$ are defined by formula (1) via the operators $Q_{r}=P_{r}^{s_{1}}\otimes...\otimes P_{r}^{s_{l}}\otimes I_{R}$, where $\{s_{1},...,s_{l}\}$ is any subset of $\{1,...,n\}$ and $R=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{l}}$. For a smaller class of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ the above result is valid with the FA-property of the states $\rho_{A_{1}}$,…, $\rho_{A_{m}}$ replaced by the weaker condition of finiteness of their entropy (Theorem 3.1 in Section 3).
The above result and its modifications are used in this article to solve the following tasks for a characteristic $f$ of composite quantum systems:
  • to obtain conditions for local lower semicontinuity and local continuity (convergence) of $f$, i.e. conditions under which the limit relations
    $\liminf_{k\to+\infty}f(\rho_{k})\geq f(\rho_{0})\quad\textrm{and}\quad\lim_{k\to+\infty}f(\rho_{k})=f(\rho_{0})$
    hold for a given sequence $\{\rho_{k}\}$ of states converging to a state $\rho_{0}$;
  • to obtain conditions under which a convex characteristic $f$ is integrable [^3] This question is not trivial, since it is unclear how to prove the measurability of many important characteristics of infinite-dimensional quantum systems. w.r.t. a Borel probability measure $\mu$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ and satisfies the Jensen inequality
    $f(\bar{\rho}(\mu))\leq\int_{\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})}f(\rho)\mu(d\rho),$ (3)
    where
@@
  • to establish a finite-dimensional approximation property for $f$ that allows us to generalize to the infinite-dimensional case the results (equalities, inequalities, etc.) proved in the finite-dimensional settings;
  • to simplify definition of $f$ for states satisfying certain regularity conditions;
  • to obtain conditions under which $f(\rho)=0$ for a separable (in different senses) state $\rho$ (when $f$ is an entanglement measure).
The article is organized as follows. In Section 2.1, we give a brief description of basic notation and describe several entropic characteristics used below. Section 2.2 is devoted to the simple sufficient condition for the FA-property mentioned before. In Section 2.3, we introduce the classes $I_{n}^{m}$ and $II_{n}^{m}$ ($1\leq m\leq n$) of functions on the set $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ characterized by the existence of faithful uniform continuity bounds with two types ($I_{n}^{m}$ and $II_{n}^{m}$) of energy constraint. These classes allow us to simplify the presentation of our main technical result (Theorem 3.1), obtained in Section 3.1. In Section 3.2, we analyse several corollaries of Theorem 3.1 concerning general properties of correlation and entanglement measures in infinite-dimensional composite quantum systems. In Section 3.3, we consider a sufficient condition for local continuity (convergence) of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ w.r.t. the strong information topology (obtained by modifying the main step of the proof of Theorem 3.1).
In Section 4 we apply the general results from Section 3 to three important characteristics of infinite-dimensional composite quantum systems which are difficult to analyse by the methods proposed before. In Section 4.1, we explore properties of the relative entropy of $\pi$-entanglement and its regularization extending the results from [6, 7]. Section 4.2 is devoted to the Rains bound (unregularized and regularized) – a function of the relative entropy distance type for which the technique used in [6, 7] cannot be applied. In Section 4.3, we obtain several results concerning properties of the conditional entanglement of mutual information in the infinite-dimensional settings.
In Section 5 we summarize the results and discuss other possible applications of the technique proposed in the article.
@@ 2.1 Basic notation
We will use Greek letters to denote operators in $\mathfrak{T}(\mathcal{H})$ and Latin letters for other operators, in particular, unbounded operators on a Hilbert space $\mathcal{H}$. The support $\mathrm{supp}\rho$ of an operator $\rho$ in $\mathfrak{T}_{+}(\mathcal{H})$ is the closed subspace spanned by the eigenvectors of $\rho$ corresponding to its positive eigenvalues. The dimension of $\mathrm{supp}\rho$ is called rank of $\rho$ and denoted by $\mathrm{rank}\rho$.
@@
Lindblad’s extension of the quantum relative entropy for any operators $\rho$ and $\sigma$ in $\mathfrak{T}_{+}(\mathcal{H})$ is defined as
$D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)=\sum_{i}\langle\varphi_{i}|\,\rho\ln\rho-\rho\ln\sigma+\sigma-\rho\,|\varphi_{i}\rangle,$ (6)
where $\{\varphi_{i}\}$ is the orthonormal basis of eigenvectors of the state $\rho$ and it is assumed that $D(0\hskip 1.0pt\|\hskip 1.0pt\sigma)=\mathrm{Tr}\sigma$ and that $D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)=+\infty$ if $\,\mathrm{supp}\rho\hskip 1.0pt$ is not contained in $\hskip 1.0pt\mathrm{supp}\hskip 1.0pt\sigma$ [10]. [^4] The support $\mathrm{supp}\rho$ of a positive trace class operator $\rho$ is the closed subspace spanned by the eigenvectors of $\rho$ corresponding to its positive eigenvalues.
@@
The quantum mutual information (QMI) of a state $\,\rho\,$ of a $n$-partite quantum system $A_{1}\ldots A_{n}$ is defined as (cf.[12, 13, 14])
$I(A_{1}\!:\ldots:\!A_{n})_{\rho}\doteq D(\rho\hskip 1.0pt\|\hskip 1.0pt\rho_{A_{1}}\otimes\cdots\otimes\rho_{A_{n}})=\sum_{s=1}^{n}S(\rho_{A_{s}})-S(\rho),$ (8)
@@
Note also that the QMI satisfies the inequalities
$I(A_{1}\!:...:\!A_{n})_{p\rho+(1-p)\sigma}\geq pI(A_{1}\!:...:\!A_{n})_{\rho}+(1-p)I(A_{1}\!:...:\!A_{n})_{\sigma}-h_{2}(p)$ (10)
and
$I(A_{1}\!:...:\!A_{n})_{p\rho+(1-p)\sigma}\leq pI(A_{1}\!:...:\!A_{n})_{\rho}+(1-p)I(A_{1}\!:...:\!A_{n})_{\sigma}+(n-1)h_{2}(p)$ (11)
for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ and $p\in(0,1)$, where $h_{2}(p)$ is the binary entropy and both sides can be equal to +\,$\infty$+ ∞. Indeed, if $\rho$ and $\sigma$ are states with finite marginal entropies then inequality (10) and (11) can be proved by using the second expression in (8), concavity of the entropy and inequality (5). The validity of (10) and (11) for arbitrary states $\rho$ and $\sigma$ can be shown by approximation using Proposition 5B in [16].
For a positive (semi-definite) operator [^5] We assume that a positive operator is a self-adjoint operator [17]. $H$ on a Hilbert space $\mathcal{H}$ and any operator $\rho$ in $\mathfrak{T}_{+}(\mathcal{H})$ we assume that
$\mathrm{Tr}H\rho=\left\{\begin{array}[]{l}\sup_{n}\mathrm{Tr}P_{n}H\rho\;\;\textrm{if}\;\;\mathrm{supp}\rho\subseteq{\rm cl}(\mathcal{D}(H))\\ +\infty\;\;\qquad\qquad\textrm{otherwise}\end{array}\right.,$ (12)
where $P_{n}$ is the spectral projector of $H$ corresponding to the interval $[0,n]$ and ${\rm cl}(\mathcal{D}(H))$ is the closure of the domain $\mathcal{D}(H)$ of $\mathcal{H}$. Then
@@
is a closed convex nonempty subset of $\mathfrak{S}(\mathcal{H})$ for any $E$ greater than the infimum of the spectrum of $H$. If $H$ is treated as Hamiltonian of a quantum system associated with the space $\mathcal{H}$ then $\mathfrak{C}_{H,E}$ is the set of states with the mean energy not exceeding $E$.
We will pay a special attention to the class of unbounded positive operators on $\mathcal{H}$ with a discrete spectrum. In Dirac’s notation any such operator $H$ can be represented as
@@
on the domain $\mathcal{D}(H)=\{\varphi\in\mathcal{H}_{\mathcal{T}}\,|\,\sum_{i=1}^{+\infty}h^{2}_{i}|\langle\tau_{i}|\varphi\rangle|^{2}<+\infty\}$, where $\mathcal{T}=\left\{\tau_{i}\right\}_{i=1}^{+\infty}$ is the orthonormal system of eigenvectors of $H$ corresponding to the nondecreasing sequence $\left\{\smash{h_{i}}\right\}_{i=1}^{+\infty}$ of eigenvalues tending to +$\infty$+ ∞ and $\mathcal{H}_{\mathcal{T}}$ is the linear span of $\mathcal{T}$.
@@
and that the maximal value of the entropy on this set is achieved at the Gibbs state $\gamma_{H}(E)\doteq e^{-\beta(E)H}/\mathrm{Tr}e^{-\beta(E)H}$, where the parameter $\beta(E)$ is determined by the equality $\mathrm{Tr}He^{-\beta(E)H}=E\mathrm{Tr}e^{-\beta(E)H}$ [3]. Condition (15) implies that $H$ is an unbounded operator with a discrete spectrum, i.e. it has the form (14). So, by the Lemma in [18] the set $\mathfrak{C}_{H,E}$ defined in (13) is compact for any $E$.
@@
If equivalent properties $(i)\textrm{-}(iii)$ hold then the positive operator $H$ in $(ii)$ can be chosen diagonizable in any basis of eigenvectors of the state $\rho$ and such that $H|\varphi_{1}\rangle=0$, where $\varphi_{1}$ is any eigenvector of $\rho$ corresponding to its maximal eigenvalue.
Proof. The implication $\rm(ii)\Rightarrow(i)$ follows from the results mentioned before the proposition. To prove $\rm(i)\Rightarrow(iii)$ it suffices to note that $\lambda^{\rho}_{i}\leq 1/i$ for all $i$.
Assume that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln i<+\infty$. Then it is easy to show the existence of an increasing sequence $\{c_{i}\}_{i=1}^{+\infty}$ tending to +$\infty$+ ∞ such that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}c_{i}\ln i<+\infty$. Then $\rm(ii)$ holds with the operator
@@
where $\{\varphi_{i}\}_{i=1}^{+\infty}$ is any basis of eigenvectors of $\rho$ corresponding to the sequence $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$ (if $\rho$ is not a faithful state then the basis $\{\varphi_{i}\}_{i=1}^{+\infty}$ is obtained by extending the basis of eigenvectors of $\rho$ in $\mathcal{H}_{\rho}\doteq\mathrm{supp}\rho$ to a basis in $\mathcal{H}$). $\Box$
@@
which is stronger than condition (15). By Lemma 1 below the condition (17) is valid if $\;\liminf\limits_{i\rightarrow+\infty}h_{i}/\ln^{2}i=+\infty$, where $\{h_{i}\}$ is the sequence of eigenvalues of $H$. By Lemma 1 in [20] condition (17) holds if and only if
@@
$F_{H^{2}}(E)\leq F_{H}(\sqrt{E})\quad\forall E\geq h^{2}_{1},$
where $h_{1}$ is the minimal eigenvalue of $H$. This inequality follows from the definitions of $F_{H^{2}}(E)$ and $F_{H}(\sqrt{E})$ due to the inequality
@@
valid for any state $\rho$ in $\mathfrak{S}(\mathcal{H})$ (with possible value +$\infty$+ ∞ in one or both sides). It is clear that $\mathcal{D}(H^{2})\subseteq\mathcal{D}(H)$. So, the last inequality is a direct corollary of the concavity of the square root, since for any state $\rho$ with finite $\mathrm{Tr}H^{2}\rho$ we have
@@
An interesting open question is whether it is possible to reverse the claim of Lemma 2.1.